Grade 12General Physics

Two bodies of masses m and 4m are released simultaneously from 64m and16m height s respectively the ratio of times taken by the bodies to reach ground is

Karthik
9 Years agoGrade 12

1 Answer

Anusha Misra
9 Years ago
So taking 2 cases, we have:
Case 1:
h= 64 m
u= 0 m/s (since the  body is "released")
a= g = 10 m/s^2
say time taken for the fall is t1 seconds
using the formula, h= ut + 1/2at^2,
64 = 1/2*10*(t1)^2
=> (t1)^2=64/5    ----------------------- (i)
Similarly, for case 2, 
say time taken for the fall is t2,
=> (t2)^2  = 16/5 ----------------------- (ii)
dividing (i) and (ii),
we get,
(t1/t2)^2=4
=> t1/t2 = 2 (since time cannot be negative)
Thus, the ratio of time taken is 2.